# FlowNav: Combining Flow Matching and Depth Priors for Efficient Navigation

### v2

@@ Abstract
Effective robot navigation in unseen environments is a challenging task that requires precise control actions at high frequencies. Recent advances have framed it as an image-goal-conditioned control problem, where the robot generates navigation actions using frontal RGB images. Current state-of-the-art methods in this area use diffusion policies to generate these control actions. Despite their promising results, these models are computationally expensive and suffer from weak perception. To address these limitations, we present FlowNav, a novel approach that uses a combination of Conditional Flow Matching (CFM) and depth priors from off-the-shelf foundation models to learn action policies for robot navigation. FlowNav is significantly more accurate at navigation and exploration than state-of-the-art methods. We validate our contributions using real robot experiments in multiple unseen environments, demonstrating improved navigation reliability and accuracy. We make the code and trained models publicly available.
@@ I Introduction
Figure 1: Comparison of sampled trajectories with goal-masked diffusion (NoMaD) and FlowNav (ours) for the same observations. Sampled trajectories are shown in blue, whereas the yellow trajectories are executed. FlowNav uses a combination of CFM and depth priors to generate collision-free actions.
@@ II-A Classical Navigation
Autonomous navigation has been a central research area in robotics, with various approaches addressing the challenge of guiding robots through complex environments [1]. Conventional methods that are based on the combination of Simultaneous Localization and Mapping (SLAM) [5, 23] and path planning techniques [24, 25] have demonstrated substantial robustness. However, these classical approaches struggle to generalize across domains and environments.
@@ II-B Learning-based Navigation
A significant body of research has also included language instructions instead of goal images [13]. SPOC [31] presents an autonomous agent trained with a shortest-path expert. Follow-up works [32, 33] introduce a zero-shot policy for navigation using reinforcement learning fine-tuning. While learning from human-comprehensible commands is a promising research area, it falls outside the scope of our work. We focus on solving the problem of visual navigation, where the robot should learn to predict actions based only on a frontal camera image.
@@ III-A Problem Formulation
Visual-navigation models try to solve the problem of goal-conditioned navigation, where a robot must either explore its surroundings or navigate to a specified image goal. We define a visual navigation policy $\pi$, which takes as input a set of observations $\prescript{i}{}{O}$ and outputs actions $\prescript{i}{}{A}$, where $i$ denotes the current image index. We define $\prescript{i}{}{o}$ as the current observation image from the camera and $T_{o}$ as the past observation horizon. The observation context includes all the images in the observation horizon $\prescript{i}{}{O}=\{\prescript{(i-T_{o})}{}{o}:\prescript{i}{}{o}\}$. The policy $\pi$ samples from the distribution $P(\prescript{i}{}{A}\mid\prescript{i}{}{O})$, where $\prescript{i}{}{A}$ is the set of predicted future actions at $i$ with a horizon of size $T_{a}$, such that $\prescript{i}{}{A}=\{\prescript{i}{}{a}:\prescript{i+T_{a}}{}{a}\}$, where $\prescript{i}{}{a}$ is a single waypoint.
@@ IV-A FlowNav Architecture
A novel addition to our architecture is the use of depth priors from off-the-shelf foundation models for monocular depth estimation. We denote the feature encoder of this model as $h$. For simplicity, we only encode the current observation $\prescript{i}{}{o}$. Although the choice of model to use for our approach is arbitrary, owing to a vast amount of research on monocular depth estimation [41, 42, 40], we use the pre-trained DinoV2 [43] encoder from the Depth-Anything-V2 model [40] because of the large amount of labeled and unlabeled data that the model was trained on. The novel student-teacher distillation framework of Depth-Anything-V2 also helps the depth model generalize to unseen data. We use the VIT-small version of the Depth-Anything-V2 model that generates a set of embeddings $h(\prescript{i}{}{o})$ $\in\mathbb{R}^{C\times D}$ with $D=384$. Further, we apply an adaptive average pooling along the channel dimension $C$ with a pooling dimension of 64 to summarize the embedding information. The flattened embedding is then processed through an MLP to generate the depth tokens $\prescript{i}{}{d}$, which are used as inputs to the next block, the state encoder.
@@ V-C Deployment
Figure 3: Example images of some of the environments that were used in the experiments. We include easy (boxes and sofas) and hard obstacles (tables and chairs). Our test environment also involves navigating through glass walls and cluttered environments.
@@ V-D Experimental Results
We report results of all models for the goal-directed navigation and in the exploratory mode. For goal-directed navigation, we report the goal-reaching success rates and the number of minor collisions per run. We consider collisions that the robot cannot recover from as major collisions, for instance, getting stuck for more than 10 seconds or frontal collisions. For the exploration experiments, the robot is tasked with following a closed path. In this case, we record the number of collisions that occur when the robot explores the path.
@@ V-D1 Baseline Comparison
Table I compares all trained models on navigation and exploration experiments using $\mathcal{K}=10$ inference steps for action generation. The results show that FlowNav outperforms NoMaD and NoMaD-with-depth in all metrics. We also show that FlowNav-without-depth has a success rate comparable to NoMaD. Moreover, NoMaD-with-depth and FlowNav have far fewer collisions than the plain models. We observe a similar trend in the exploration experiments, where FlowNav model has just a single collision per run, which is a significant improvement over the plain models. This reinstates our hypothesis that including depth priors during model training helps the model learn better action policies.
@@ V-D2 Qualitative Results
Fig. 4 illustrates trajectories generated for navigation and exploration by FlowNav in various environments. The images show a set of generated actions (in blue and yellow). The robot executes the action in yellow, which is selected at random. We observe that FlowNav accurately learns contextually aware action distributions for exploration and navigation.
Figure 5: Comparison of generated trajectories with NoMaD-with-depth and FlowNav for $\mathcal{K}=2,5$ and 10 inference steps. FlowNav produces valid trajectories with fewer steps, showcasing the efficiency of flow matching over diffusion policies.
@@ V-D4 Inference Steps Ablation
To examine how the number of inference steps $\mathcal{K}$ impacts model performance, we conduct experiments in simplified environments with soft turns and no obstacles. Since we have already demonstrated that depth-conditioned models perform better, we limit our comparison to these models. We evaluate navigation and exploration tasks while varying $\mathcal{K}$. For navigation, we again track success rates and the number of minor collisions per run, as in the previous experiment. For exploration, as we use environments without obstacles, we only track if exploration was successful (no collisions). The results of these experiments are presented in Tables II and III. We observe that NoMaD-with-depth performs poorly with a small $\mathcal{K}$ but achieves a high success and a low collision rate at $\mathcal{K}=10$. In contrast, FlowNav performs well even at $\mathcal{K}=2$ with a $75\%$ success rate, and continues to improve as $\mathcal{K}$ increases. We observe a similar trend in the exploration experiments, where FlowNav performs well at $\mathcal{K}=2$, while NoMaD-with-depth struggles to explore the environment due to frequent collisions. These findings highlight that flow matching policies remain effective with fewer inference steps, thus reducing trajectory generation time. Fig. 5 visually represents our findings.
@@ V-D5 Inference Time
Table IV shows action inference times with different numbers of inference steps $\mathcal{K}$. As seen in Table I, NoMaD-with-depth outperforms NoMaD. We also show in Tables II, III and Fig. 5 that FlowNav at $\mathcal{K}=2$ is comparable to NoMaD-with-depth at $\mathcal{K}=10$. Thus, FlowNav at $\mathcal{K}=2$ produces better actions compared to NoMaD at $\mathcal{K}=10$, while being faster (about 3x) to generate these actions, as reported in Table IV.
TABLE IV: Inference times (in ms) for different no. of inference steps
@@ VI Conclusions
In this paper, we present FlowNav as a novel and image-conditioned action policy that is trained with Conditional Flow Matching (CFM) and additionally integrates depth priors. The improved efficiency obtained through the usage of CFM and the addition of the depth priors leads to a better performance in both navigation and exploration compared to the state-of-the-art approach NoMaD. In particular, FlowNav demonstrates substantially improved obstacle avoidance capabilities. We validate this through extensive real-robot experiments. For future work, we would like to focus on using off-policy reinforcement learning to learn an optimal trajectory for goal-based exploration. Moreover, we plan to substitute the topological map with implicit representations learned with the same transformer backbone.

### v3

@@ Abstract
Effective robot navigation in unseen environments is a challenging task that requires precise control actions at high frequencies. Recent advances have framed it as an image-goal-conditioned control problem, where the robot generates navigation actions using frontal RGB images. Current state-of-the-art methods in this area use diffusion policies to generate these control actions. Despite their promising results, these models are computationally expensive and suffer from weak perception. To address these limitations, we present FlowNav, a novel approach that uses a combination of Conditional Flow Matching (CFM) and depth priors from off-the-shelf foundation models to learn action policies for robot navigation. FlowNav is significantly more accurate and faster at navigation and exploration than state-of-the-art methods. We validate our contributions using real robot experiments in multiple environments, demonstrating improved navigation reliability and accuracy. Code and trained models are publicly available.
@@ I Introduction
Figure 1: Comparison of sampled trajectories with goal-masked diffusion (NoMaD) and FlowNav (ours). Sampled trajectories are in blue; yellow trajectories are executed. FlowNav uses a combination of CFM and depth priors to generate collision-free actions.
@@ II-A Classical Navigation
Autonomous navigation has been a central research area in robotics, with various approaches addressing the challenge of guiding robots through complex environments [1]. Conventional methods that are based on the combination of Simultaneous Localization and Mapping (SLAM) [5, 23] and path planning techniques [24, 25] have demonstrated substantial robustness. However, these classical approaches often depend on hyperparameters that are environment specific, struggling to generalize across domains and environments.
@@ II-B Learning-based Navigation
A significant body of research has also included language instructions instead of goal images [13]. SPOC [31] presents an autonomous agent trained with a shortest-path expert. Follow-up works [32, 33] introduce a zero-shot policy for navigation using reinforcement learning fine-tuning. While learning from human-comprehensible commands is a promising research area, it is considered orthogonal work and falls outside our scope. We focus on solving the problem of visual navigation, where the robot should learn to predict actions based only on a frontal camera image.
@@ III-A Problem Formulation
Visual-navigation models try to solve the problem of goal-conditioned navigation, where a robot either explores its surroundings or navigates to a specified image goal. We define a visual navigation policy $\pi$, which takes as input a set of observations $\prescript{i}{}{O}$ and outputs actions $\prescript{i}{}{A}$, where $i$ denotes the current image index. We define $\prescript{i}{}{o}$ as the current observation image from the camera and $T_{o}$ as the past observation horizon. The observation context includes all the images in the observation horizon $\prescript{i}{}{O}=\{\prescript{(i-T_{o})}{}{o}:\prescript{i}{}{o}\}$. The policy $\pi$ samples from the distribution $P(\prescript{i}{}{A}\mid\prescript{i}{}{O})$, where $\prescript{i}{}{A}$ is the set of predicted future actions at $i$ with a horizon of size $T_{a}$, such that $\prescript{i}{}{A}=\{\prescript{i}{}{a}:\prescript{i+T_{a}}{}{a}\}$, where $\prescript{i}{}{a}$ is a single waypoint.
@@ IV-A FlowNav Architecture
A novel addition to our architecture is the use of depth priors from off-the-shelf foundation models for monocular depth estimation. We denote the feature encoder of this model as $h$. For simplicity, we only encode the current observation $\prescript{i}{}{o}$. Although the choice of model to use for our approach is arbitrary, owing to a vast amount of research on monocular depth estimation [41, 42, 40], we use the pre-trained DinoV2 [43] encoder from the Depth-Anything-V2 model [40] because of the large amount of labeled and unlabeled data that the model was trained on. We use the VIT-small version of the Depth-Anything-V2 model that generates a set of embeddings $h(\prescript{i}{}{o})$ $\in\mathbb{R}^{C\times D}$ with $D=384$. Further, we apply an adaptive average pooling along the channel dimension $C$ with a pooling dimension of 64 to summarize the embedding information. The flattened embedding is then processed through an MLP to generate the depth tokens $\prescript{i}{}{d}$, which are used as inputs to the next block, the state encoder.
@@ V-C Deployment
Figure 3: Examples of environments. We include easy (boxes and sofas) and hard obstacles (tables and chairs). Our test environment also involves navigating through glass walls and cluttered environments.
@@ V-D Experimental Results
We report results of all models for the goal-directed navigation and in the exploratory mode. For goal-directed navigation, we report the goal-reaching success rates and the number of minor collisions per run. We consider collisions that the robot cannot recover from as major collisions, for instance, getting stuck for more than 10 seconds or frontal collisions. For the exploration experiments, the robot follows a closed path. In this case, we record the number of collisions that occur when the robot explores the path.
@@ V-D1 Baseline Comparison
Since NoMaD outperforms [10] the previous models GNM [11] and ViNT [12], we compare to NoMaD as the sole baseline. Table I compares all trained models on navigation and exploration experiments using $\mathcal{K}=10$ inference steps for action generation. The results show that FlowNav outperforms NoMaD and NoMaD-with-depth in all metrics. We also show that FlowNav-without-depth has a success rate comparable to NoMaD. Moreover, NoMaD-with-depth and FlowNav have far fewer collisions than the plain models. We observe a similar trend in the exploration experiments, where FlowNav model has just a single collision per run, which is a significant improvement over the plain models. This reinstates our hypothesis that including depth priors during model training helps the model learn better action policies.
@@ V-D2 Qualitative Results
Fig. 4 illustrates trajectories generated for navigation and exploration by FlowNav in various environments. The images show a set of generated actions (in blue and yellow). The robot executes the action in yellow, which is selected at random. FlowNav accurately learns contextually aware actions for exploration and navigation.
Figure 5: Comparison of generated trajectories with NoMaD-with-depth and FlowNav for $\mathcal{K}=2,5$ and 10 inference steps. FlowNav produces valid trajectories with fewer steps, showing the effectiveness of flow matching.
@@ V-D4 Inference Steps Ablation
[…]
@@ V-D5 Inference Time
Table IV shows action runtime with different numbers of inference steps $\mathcal{K}$. We show in Tables II, III and Fig. 5 that FlowNav at $\mathcal{K}=2$ is comparable to NoMaD-with-depth at $\mathcal{K}=10$. Thus, even though FlowNav has comparable runtime to NoMaD for the same $\mathcal{K}$, it requires less steps to beat the performance of NoMaD.
TABLE IV: Inference times (in ms) for different no. of steps
@@ VI Conclusions
In this paper, we present FlowNav as a novel and image-conditioned action policy that is trained with Conditional Flow Matching (CFM) and additionally integrates depth priors. The improved efficiency obtained through the usage of CFM and the addition of the depth priors leads to a better performance and faster runtimes in both navigation and exploration compared to the state-of-the-art approach NoMaD. We validate this through extensive real-robot experiments. Our results show that FlowNav consistently outperforms prior approaches, setting a new standard for robust and efficient navigation. For future work, we would like to focus on using off-policy reinforcement learning to learn an optimal trajectory for goal-based exploration. We plan to substitute the topological map with implicit representations learned with the same transformer backbone.